# Multiplier modules of Hilbert C*-modules revisited

### v2

@@ 3. On modular operators and functionals
The aim of this section is the investigation of the Banach algebras of all bounded module maps ${\rm End}_{A}(X)$ and ${\rm End}_{M(A)}(M(X))$ and their interrelations, as well as the sets of all bounded module maps $X^{\prime}$ and $M(X)^{\prime}$ over pairs of (full) Hilbert $A$-modules $X$ and their multiplier modules $M(X)$ over $M(A)$. To get non-trivial examples we need examples of C*-algebras $A$ such that their multiplier algebras $M(A)$ are strictly smaller than their left/right multiplier algebras. By [8, Cor. 4.18] either $M(A)=LM(A)$ and $M(A)=QM(A)$ at the same time, or $LM(A)$ is strictly larger than $M(A)$ and $QM(A)$ is strictly larger than $LM(A)$. For theory and examples see the existing literature on different types of multiplier algebras and local multiplier algebras, e.g. [7]. We give a simple example following [32, pp. 165-166].
Recall the C*-algebras $c_{0}$, $c$ and $l_{\infty}$ of all complex-valued sequences converging to zero, converging at all and being bounded in norm, respectively. Change the target C*-algebra $\mathbb{C}$ to the C*-algebra of all two-by-two valued matrices $M_{2}({\mathbb{C}})$. Consider the C*-algebra of all $M_{2}({\mathbb{C}})$-valued sequences with the sequence in the upper left corner converging at all and with the sequences derived from the other three positions being sequences converging to zero. We write $A$ as a symbol
@@
Restricting a bounded $M(A)$-linear operator on $M(X)$ on the domain from $M(X)$ to $X\subseteq M(X)$ one obtains a bounded $A$-linear operator on $X$. The norm is preserved since $X$ is strictly dense in $M(X)$, there does not exist any non-trivial bounded module operator on $M(X)$ vanishing on $X\subseteq M(X)$ and the norm is preserved on the subalgebra ${\rm End}_{(M(A)}(M(X))^{*}\equiv{\rm End}_{A}(X)^{*}$ by [5, Thm. 2.3]. So, the restriction of a non-adjointable operator is a non-adjointable operator, again. In case, the left multiplier algebra $LM({\rm K}_{A}(X))={\rm End}_{A}(X)$ is larger than the multiplier algebra $M({\rm K}_{A}(X))={\rm End}_{A}(X)^{*}$ (see example above) not all elements of $LM({\rm K}_{A}(X))\setminus M({\rm K}_{A}(X))$ might be extendable to bounded module operators on $M(X)$: Indeed, if $A=X$ is a C*-algebra with $LM(A)\supset M(A)$ then $M(X)=M(A)$ and all elements of $LM(A)\setminus M(A)$ cannot be continued from $X\subset M(X)$ to $M(X)$, cf. [33, Thm. 1.5, 1.6], [8, Cor. 4.18]. ∎
@@
We demonstrate by example that the choice of the $A$-valued inner product on Hilbert $A$-modules $X$ within the class of $A$-valued inner products on $X$ inducing equivalent norms on $X$ may lead to other unitarily non-equivalent multiplier $M(A)$-modules. Return to the example at the beginning of the present section. If the C*-algebra $A$ defined there is equipped with the standard $A$-valued inner product as a Hilbert $A$-module $X$ then $M(X)=M(A)$. Now, modify this $A$-valued inner product setting .,.$\langle.,.\rangle_{1}:=\langle T(.),T(.)\rangle_{A}$ for
@@
where 00 is the zero sequence and $1$ is the identity sequence. Clearly, $T$ is a non-adjointable invertible bounded $A$-linear operator on $A$, and .,.$\langle.,.\rangle_{1}$ is an $A$-valued inner product on $A$ inducing an equivalent Hilbert module norm on $A=X$. A simple calculation for elements of $M(A)$ inside the C*-valued inner product .,.$\langle.,.\rangle_{1}$ yields
@@
Consequently, the $A$-valued inner product .,.$\langle.,.\rangle_{1}$ cannot be extended to $M(A)$, and $M(A)$ is not the multiplier module of the Hilbert $A$-module $X_{1}=\{A,\langle.,.\rangle_{1}\}$. So, the Banach $A$-module $A$ of the concrete example does not determine its multiplier module alone, one has to take into account the particular $A$-valued inner product on it. To calculate the multiplier module of $X_{1}$ one can use the von Neumann algebra of all bounded $M_{2}({\mathbb{C}})$-valued sequences as an environment. Then
@@
Restricting $f\in M(X)^{\prime}$ to $X\subseteq M(X)$ we obtain a bounded $A$-linear functional of $X^{\prime}$. The norm is preserved, since $X$ ist strictly dense in $M(X)$ and there does not exist any non-trivial bounded $M(A)$-linear functional on $M(X)$ vanishing on $X\subseteq M(X)$. The example in the beginning of the present section can be read as follows: Let $A$ be a C*-algebra such that $LMA)\supset M(A)$. Setting $A=X$ we get $X^{\prime}=LM(A)$ and $M(X)^{\prime}=M(A)^{\prime}=M(A)$ since $M(X)=M(X)^{\prime}$ is selfdual. So $X^{\prime}\supset M(X)^{\prime}$ and a functional in $LM(A)\setminus M(A)$ cannot be continued from $X\subseteq M(X)$ to $M(X)$. ∎
Consequently, there does not exist any general Hahn-Banach type theorem for bounded C*-linear functionals for pairs $(X,M(X))$ of full Hilbert C*-modules $X$ and their multiplier modules $M(X)$, cf. [20, 21]. However, if a bounded $A$-linear functional from a full Hilbert $A$-module to the C*-algebra $A$ admits a continuation to a bounded $M(A)$-linear functional from its multiplier module to $M(A)$ of the same norm value then it is unique.
@@
The example in the beginning of the present section shows: Let $A$ be a C*-algebra such that $LMA)\supset M(A)$. Setting $A=X$ we get ${\rm End}_{A}(X,X^{\prime})=QM({\rm K}_{A}(X))$ and ${\rm End}_{M(A)}(M(X),M(X)^{\prime})=QM({\rm K}_{M(A)}(M(X))$, cf. [33, Thm. 1.6]. Since $QM({\rm K}_{A}(X))\not=LM({\rm K}_{A}(X))$, but $QM({\rm K}_{M(A)}(M(X)))=M({\rm K}_{M(A)}(M(X)))=M(M(A))=M(A)$ by [8, Cor. 4.18], the assertion is demonstrated. ∎
@@
Further research directions include the open problem whether multiplier modules are C*-reflexive, more precisely, $M(A)$-reflexive. If not always, the exceptions have to be characterized. The background of this problem is a result by W. L. Paschke that the C*-valued inner product on a full Hilbert C*-module can be always continued to its C*-reflexive C*-bidual Banach C*-module preserving the isometric embedding of the former into the latter, cf. [38].

### v3

@@ 3. On modular operators and functionals
The aim of this section is the investigation of the Banach algebras of all bounded module maps ${\rm End}_{A}(X)$ and ${\rm End}_{M(A)}(M(X))$ and their interrelations, as well as the sets of all bounded module maps $X^{\prime}$ and $M(X)^{\prime}$ over pairs of (full) Hilbert $A$-modules $X$ and their multiplier modules $M(X)$ over $M(A)$. To get non-trivial examples we need examples of C*-algebras $A$ such that their multiplier algebras $M(A)$ are strictly smaller than their left/right multiplier algebras. By [9, Cor. 4.18] either $M(A)=LM(A)$ and $M(A)=QM(A)$ at the same time, or $LM(A)$ is strictly larger than $M(A)$ and $QM(A)$ is strictly larger than $LM(A)$. For theory and examples see the existing literature on different types of multiplier algebras and local multiplier algebras, e.g. [8]. We give a simple example following [35, pp. 165-166].
Recall the C*-algebras $_{0}, cc$ and $l_{\infty}$ of all complex-valued sequences converging to zero, converging at all and being bounded in norm, respectively. Change the target C*-algebra $\mathbb{C}$ to the C*-algebra of all two-by-two valued matrices $M_{2}({\mathbb{C}})$. Consider the C*-algebra of all $M_{2}({\mathbb{C}})$-valued sequences with the sequence in the upper left corner converging at all and with the sequences derived from the other three positions being sequences converging to zero. We write $A$ as a symbol
@@
Restricting a bounded $M(A)$-linear operator on $M(X)$ on the domain from $M(X)$ to $X\subseteq M(X)$ one obtains a bounded $A$-linear operator on $X$. The norm is preserved since $X$ is strictly dense in $M(X)$, there does not exist any non-trivial bounded module operator on $M(X)$ vanishing on $X\subseteq M(X)$ and the norm is preserved on the subalgebra ${\rm End}_{(M(A)}(M(X))^{*}\equiv{\rm End}_{A}(X)^{*}$ by [5, Thm. 2.3]. So, the restriction of a non-adjointable operator is a non-adjointable operator, again. In case, the left multiplier algebra $LM({\rm K}_{A}(X))={\rm End}_{A}(X)$ is larger than the multiplier algebra $M({\rm K}_{A}(X))={\rm End}_{A}(X)^{*}$ (see example above) not all elements of $LM({\rm K}_{A}(X))\setminus M({\rm K}_{A}(X))$ are extendable to bounded module operators on $M(X)$: Indeed, if $A=X$ is a C*-algebra with $LM(A)\supset M(A)$ then $M(X)=M(A)$ and all elements of $LM(A)\setminus M(A)$ cannot be continued from $X\subset M(X)$ to $M(X)$, cf. [36, Thm. 1.5, 1.6], [9, Cor. 4.18]. ∎
@@
We demonstrate by example that the choice of the $A$-valued inner product on Hilbert $A$-modules $X$ within the class of $A$-valued inner products on $X$ inducing equivalent norms on $X$ may lead to other unitarily non-equivalent multiplier $M(A)$-modules. Return to the example at the beginning of the present section. If the C*-algebra $A$ defined there is equipped with the standard $A$-valued inner product as a Hilbert $A$-module $X$ then $M(X)=M(A)$. Now, modify this $A$-valued inner product setting $\langle.,.\rangle_{1}:=\langle T(.),T(.)\rangle_{A}$ for
@@
where 0 is the zero sequence and $1$ is the identity sequence. Clearly, $T$ is a non-adjointable invertible bounded $A$-linear operator on $A$, and $\langle.,.\rangle_{1}$ is an $A$-valued inner product on $A$ inducing an equivalent Hilbert module norm on $A=X$. A simple calculation for elements of $M(A)$ inside the C*-valued inner product $\langle.,.\rangle_{1}$ yields
@@
Consequently, the $A$-valued inner product $\langle.,.\rangle_{1}$ cannot be extended to $M(A)$, and $M(A)$ is not the multiplier module of the Hilbert $A$-module $X_{1}=\{A,\langle.,.\rangle_{1}\}$. So, the Banach $A$-module AA of the concrete example does not determine its multiplier module alone, one has to take into account the particular $A$-valued inner product on it. To calculate the multiplier module of $X_{1}$ one can use the von Neumann algebra of all bounded $M_{2}({\mathbb{C}})$-valued sequences as an environment. Then
@@
Restricting $f\in M(X)^{\prime}$ to $X\subseteq M(X)$ we obtain a bounded $A$-linear functional of $X^{\prime}$. The norm is preserved, since $X$ ist strictly dense in $M(X)$ and there does not exist any non-trivial bounded $M(A)$-linear functional on $M(X)$ vanishing on $X\subseteq M(X)$. The example in the beginning of the present section can be read as follows: Let $A$ be a C*-algebra such that $LMA)\supset M(A)$. Setting $A=X$ we get $X^{\prime}=LM(A)$ and $M(X)^{\prime}=M(A)^{\prime}=M(A)$ since $M(X)=M(X)^{\prime}$ is selfdual. So $X^{\prime}\supset M(X)^{\prime}$ and an $A$-valued bounded functional in $LM(A)\setminus M(A)$ cannot be continued from $X\subseteq M(X)$ to $M(X)$. ∎
Consequently, there does not exist any general Hahn-Banach type theorem for bounded C*-linear functionals for pairs $(X,M(X))$ of full Hilbert C*-modules $X$ and their multiplier modules $M(X)$, cf. [22, 23]. However, if a bounded $A$-linear functional from a full Hilbert $A$-module to the C*-algebra $A$ admits a continuation to a bounded $M(A)$-linear functional from its multiplier module to $M(A)$ of the same norm value, then it is unique.
@@
The example in the beginning of the present section shows: Let $A$ be a C*-algebra such that $LMA)\supset M(A)$. Setting $A=X$ we get ${\rm End}_{A}(X,X^{\prime})=QM({\rm K}_{A}(X))$ and ${\rm End}_{M(A)}(M(X),M(X)^{\prime})=QM({\rm K}_{M(A)}(M(X))$, cf. [36, Thm. 1.6]. Since $QM({\rm K}_{A}(X))\not=LM({\rm K}_{A}(X))$, but $QM({\rm K}_{M(A)}(M(X)))=M({\rm K}_{M(A)}(M(X)))=M(M(A))=M(A)$ by [9, Cor. 4.18], the assertion is demonstrated. ∎
@@
Remark 3.5.
One natural question is whether multiplier modules might be C*-reflexive with respect to certain C*-algebras $B$ with $A\subseteq B\subseteq M(A)$, or not. The background are Definition 2.1, and a result by W. L. Paschke that the C*-valued inner product on a full Hilbert C*-module can be always continued to its C*-reflexive C*-bidual Banach C*-module preserving the isometric embedding of the former into the latter, cf. [41]. One could hope to find intrinsic topological or order-algebraical characterizations of some C*-reflexive Hilbert C*-modules, cf. the $l_{2}(K(l_{2}))$-example above in the context of [40]. In general, C*-reflexivity of Hilbert C*-modules highly depends on both the inner structures of the C*-algebra of coefficients and of the Hilbert C*-module under consideration.
To test this point of view select a non-unital C*-algebra $A$ with $M(A)=LM(A)$. First, if one considers $A$-reflexivity of $X=A$ with the standard $A$-valued inner product then $M_{A}(X)=M(A)$, $X_{A}^{\prime}=M(A)$ and $X_{A}^{\prime\prime}=A$, so $M_{A}(X)$ is not $A$-reflexive. Furthermore, considering $X=A$ as a Hilbert $M(A)$-module, $M_{M(A)}(X)=A$, $X_{M(A)}^{\prime}=M(A)$ and $X_{M(A)}^{\prime\prime}=M(A)$, so it is not $M(A)$-reflexive, either. The deeper reason could be that condition (iv) of Definition 2.1 might be violated for the C*-reflexive Hilbert C*-module extension of $X$.